OpenStudy (anonymous):

Solve for x. (2x^2)(e^x) = 3xe^x

OpenStudy (anonymous):

Use LN

OpenStudy (anonymous):

\[2x^2e^x=3xe^x\iff 2x^2=3x\]

OpenStudy (anonymous):

since \(e^x\) is never zero, you can divide both sides of your equation by \(e^x\) \(2x^2=3x\) should be easy to solve right?

OpenStudy (anonymous):

Yes! Thanks!!